OpenStudy (anonymous):

graph the line containig the given pair of points and find the slope (0,1),(-1,-2)

OpenStudy (anonymous):

m=(-2-1)/(-1-0) m=3 y-1=3(x-0) y=3x+1

OpenStudy (anonymous):

is the slope 3. how would i graph this????

OpenStudy (anonymous):

over one unit, up three units

OpenStudy (anonymous):

you already have two points to plot. connect the dots and keep going.